An individual has $30,000 invested in a stock with a beta of 0.6 and another $40,000 invested in a stock with a beta of 1.5. If these are the only two investments in his portfolio, what is his portfolio beta?

a.

1.11

b.

0.80

c.

4.03

d.

1.05

e.

0.99
Answer #1

portfolio beta = weighted average beta of individual stocks

total portfolio value = 30000 + 40000 = 70000

beta = (30000/70000) * 0.6 + (40000/70000) * 1.5

= 1.11

hence choose a)